OpenStudy (anonymous):

Find the common ratio of the geometric sequence whose first term is 3 and sixth is 96.

OpenStudy (anonymous):

Help/

OpenStudy (anonymous):

ummmm...... i don't really understand the question :P what do you mean, the common ratio??? like, the nth term?? we can't do THAT unless we have all the terms.

OpenStudy (anonymous):

first term=a=3 6th term=T6=ar^5 subistitute a in the 2nd equation and solve

OpenStudy (anonymous):

3r^5=96 r^5=96/3

OpenStudy (anonymous):

5th root to both sides and u get the ratio ;

OpenStudy (anonymous):

so r = 2

OpenStudy (anonymous):

what does ar means?

OpenStudy (anonymous):

a=First term . r=common ratio . where the nth term in G.s = ar^(n-1)

OpenStudy (anonymous):

Ohhhhhh..Thanks! :D

OpenStudy (anonymous):

Yrw